What is the major organic product of the following reaction? 1. R BH 2. H2O2, HO = 3.807.40. O OH Multiple Choice Oi
Answer #1

C4 - =CH R2-BH on CH₃-CH=CH-BR2 HOR, ReR CH=CH-CH₂ H₂O2,01 RR Tt 0-07 the-Heste ROOR R - 2o-oh A BOR o-CH=CH-CH₂ O-CH=CH-eliz